OpenStudy (anonymous):

How to solve square root of 3 sin theta=cos theta

OpenStudy (jdoe0001):

\(\bf \sqrt{3sin(\theta)}=cos(\theta)\quad ?\)

OpenStudy (anonymous):

probably

OpenStudy (anonymous):

what would that be?

OpenStudy (anonymous):

you still there?

OpenStudy (jdoe0001):

yes :)

OpenStudy (jdoe0001):

\(\bf \sqrt{3sin(\theta)}=cos(\theta)\quad \implies 3sin(\theta)=cos^2(x)\\ \quad \\ \textit{using the trig identity of }cos^2(x)=1-sin^2(x)\\ \quad \\ 3sin(\theta)=1-sin^2(x)\implies 0=-sin^2(x)-3sin(\theta)+1\)

OpenStudy (jdoe0001):

I don't see it factoring simple enough, so you may end up using the quadratic formula

OpenStudy (anonymous):

thanks man

OpenStudy (jdoe0001):

yw